# Hamiltonian Engineering of collective XYZ spin models in an optical cavity: From one-axis twisting to two-axis counter twisting models

### v1

@@ Abstract
Quantum simulation using synthetic quantum systems offers unique opportunities to explore open questions in many-body physics and a path for the generation of useful entangled states. Nevertheless, so far many quantum simulators have been fundamentally limited in the models they can mimic. Here, we are able to realize an all-to-all interaction with arbitrary quadratic Hamiltonian or effectively an infinite range tunable Heisenberg XYZ model. This is accomplished by engineering cavity-mediated four-photon interactions between 700 rubidium atoms in which we harness a pair of momentum states as the effective pseudo spin or qubit degree of freedom. Using this capability we realize for the first time the so-called two-axis counter-twisting model, an iconic XYZ collective spin model that can generate spin-squeezed states that saturate the Heisenberg limit bound. The versatility of our platform to include more than two relevant momentum states, combined with the flexibility of the simulated Hamiltonians by adding cavity tones opens rich opportunities for quantum simulation and quantum sensing in matter-wave interferometers and other quantum sensors such as optical clocks and magnetometers.
Introduction
@@
The ability to create and control different many-body interactions is key for entanglement generation, optimization, quantum sensing, and quantum simulation. Long-range interacting systems have recently shown much promise for the engineering of Hamiltonians that generate interesting correlations that propagate across the system. Several experimental platforms are making rapid progress, including Rydberg atoms Browaeys2020 ; Weidemuller2021Floquet ; Wu2021 , polar molecules Bohn2017 ; Ye2023Itinerant , trapped ions Monroe2017TimeCrystal ; Monroe2021 , cavity QED systems Mivehvar2021 and defect centers in solids Lukin2020RobustEngineering ; Munns2023 . Short range contact interactions in ultra-cold atomic systems are another promising approach Bloch2008 ; Schafer2020 ; Gross2017 , however, it remains an open challenge to reach the sufficiently low-temperatures.
@@
So far, most efforts to engineer Hamiltonians have been limited to XXZ spin models or models that feature both exchange and Ising interactions. Common to these models is the fact that the total magnetization of the spin ensemble is preserved. However, limited progress has been achieved in engineering more general spin models, such as XYZ models, which can break both SU(2), and U(1) symmetries and lead to more general ground-state and out-of-equilibrium many-body behaviors. A few exceptions include experiments in Rydberg atoms using two-color dressing Gross2023RydbergXYZ with dynamics limited to pairs of atoms, or Floquet engineering in disordered arrays Weidemuller2021Floquet .
@@
XYZ Hamiltonians can in fact provide important gains in the context of quantum sensing and metrology applications. For instance, the so called two-axis counter-twisting (TACT) KitagawaUedaOAT model is one type of XYZ model well known for its capability to generate entanglement exponentially fast all the way to the Heisenberg bound. It was proposed more than 30 years ago shirasaki1990squeezing ; KitagawaUedaOAT together with a simpler version, the so-called one-axis twisting (OAT) featuring only Ising interactions. So far, only the OAT has been realized in experiments including trapped ions Monroe2000EntanglingFourIons ; Blatt2008IonOAT , Bose-Einstein Condensates OberthalerSqueezedBECInt ; Treutlein2010Squeezing , atomic cavity-QED leroux2010implementation ; VladanEntangledClockLifetime2010 ; norcia2018cavity ; Schleier-Smith2020GapProtection , superconducting qubits Zhu2019SCqubits and optical interferometers AndrianovLeuchs2023KerrSqueezing . Experiments have approximated TACT dynamics locally on the Bloch sphere by combining OAT with a transverse drive (the so-called Lipkin-Meshkov-Glick (LMG) model) Oberthaler2015TaT ; Jessen2007CsTaT ; li2023improving or equivalent approximations in spin-nematic dynamics in higher spin systems ChapmanSqueezing2012 ; You2017TwinFock . However, there has been no demonstration of genuine TACT in any platforms due to the challenge of how to realize this more complex non-linear Hamiltonian.
Here we experimentally show that photon-mediated interactions between atoms inside an optical cavity can be tuned to realize a tunable all-to-all Heisenberg XYZ Hamiltonian, including the TACT model, with the added benefit of not requiring us to prepare extremely low entropy quantum degenerate gases. While there have been prior theory proposals to realize TACT models in atom-cavity systems using four laser tones MonikaSorenson2017NJP ; zhang2017cavity , here we experimentally realize the model using only two dressing lasers. We demonstrate the tunability of the XYZ Hamiltonian by explicitly mapping out the evolution of Bloch vectors on the Bloch sphere at the mean-field level and witness the TACT dynamics for the first time.
@@
The pseudo-spin system here is novel, consisting of two momentum states of atoms freely-falling inside the cavity, making our results of great interest for Bragg matter-wave interferometers LeePRLBragg1995 that are important for both inertial navigation and fundamental science such as searches for dark matter and dark energy, detection of gravitational waves, and determination of the fine structure constant MullerFineStructureAI ; GuellatiKhelifaFineStructure2020 . This approach can also be straightforwardly applied to systems with additional internal levels, making it ideal for developing next-generation quantum-enhanced sensors for technology and exploring a broad range of science from atomic clocks VuleticSqueezedOpticalClock2020 ; robinson2022direct and magnetometers PolzikSqueezedMagnetometer2010 ; MitchellSubSQLMagnetometry2010 ; OberthalerSqueezedMagnetometer2014 to geodesy LudlowGeodesy2018 .
@@
Figure 1: Experimental overview. (A) Illustration of the microscopic momentum pair raising process described by the Hamiltonian $\hat{J}_{+}\hat{J}_{+}$, using two momentum states $p_{0}\pm\hbar k$ as a pseudo-spin 1/2 degree of freedom. Initially, two atoms $i$ and $j$ are in the same momentum state along the cavity axis. Dressing lasers are applied to the cavity (red and blue arrows) that allow atom $i$ to absorb a dressing laser photon and emit a photon (squiggly line) into the cavity such that the net photon recoil flips its momentum state by $2\hbar k$. The emitted photon is absorbed by atom $j$, also flipping its momentum state by $2\hbar k$. There also exists the separate momentum exchange process $\hat{J}_{+}\hat{J}_{-}$, where atoms initially in opposite momentum states flip their momentum states by emitting and absorbing photons explored in Thompson2023MomentumExchange . (B) Frequency diagram of the applied dressing lasers with frequencies $\omega_{1,2}$ and coherent state amplitudes $\alpha_{1,2}$. The emitted photons (squiggly line) are Doppler shifted by $\omega_{z}$ from the dressing laser frequencies and detuned by $\Delta_{c}$ from the cavity resonance frequency $\omega_{c}$. The cavity is far detuned from the atomic transition frequency $\omega_{a}$. (C) Representation of the emission/absorption processes described in (A) but depicted in terms of the atomic energy versus momentum for atoms $i$ and $j$. The reverse process is also allowed, giving rise to a collective lowering operator described by $\hat{J}_{-}\hat{J}_{-}$. (D) Dynamics induced by two-axis counter twisting in the form of a Hamiltonian $\hat{J}_{x}^{2}-\hat{J}_{z}^{2}$ is represented on a Bloch sphere with north and south poles defined by |$\ket{p_{0}\pm\hbar k}$. Top right shows the local circular flows around the stable point at $-\hat{x}$. The local flows for unstable saddle points at $\hat{y}$ with exponential squeezing and anti-squeezing along $\hat{x}\pm\hat{z}$ are shown in bottom right. (E) Observation of the four-photon resonance that generates pair raising and lowering processes. We scan the dressing laser frequency difference to vary the detuning $\delta$. With equal dressing laser amplitudes corresponding to realizing the $\hat{J}_{x}^{2}$ Hamiltonian when $\delta=0$, we see clear resonances in the observed change of the Bloch vector’s azimuthal angle $\phi$ for a Bloch vector prepared near the south pole ($\theta=\pi/4$, red data points) and near the north pole ($\theta=3\pi/4$, blue data points). All error bars reported are $1\sigma$ uncertainties. Simulation results are shown in solid lines with shaded area allowing for 5% uncertainty in interaction strength. See Methods for details.
Experimental setup
@@
In the experiment, ${}^{87}$b atoms are laser-cooled inside a vertically-oriented two-mirror standing wave cavity, see Fig. 1(A) and Thompson2023MomentumExchange ; Thompson2022SAI . A repulsive intra-cavity doughnut dipole trap confines the atoms radially, but allows the atoms to fall along the cavity axis. To prepare atoms in a well-defined momentum state, a pair of laser beams are injected into the cavity to drive velocity-dependent two-photon Raman transitions between ground hyperfine states |$\ket{F=1,m_{F}=0}$ and |$\ket{F=2,m_{F}=0}$. After removing the unselected atoms with a resonant laser push beam, successive microwave pulses are applied to prepare the internal states of the selected atoms in the ground hyperfine state |$\ket{F=2,m_{F}=2}$ (see Methods and Thompson2023MomentumExchange ).
@@
With about 700 atoms centered at momentum $p_{0}-\hbar k$, another pair of laser beams is injected along the cavity axis to drive two-photon Bragg transitions connecting the two momentum states |$\ket{p_{0}\pm\hbar k}$ which defines a two-level spin-1/2 system Thompson2022SAI ; Thompson2023MomentumExchange . Here the average momentum is $p_{0}$, $\hbar$ is the reduced Plank constant, and $k=2\pi/\lambda$ where $\lambda$ is the wavelength of the Bragg laser beams. Due to finite momentum spread, one can consider momentum wave packets centered at these two momentum states as considered in detail previously Thompson2023MomentumExchange . Ignoring the finite momentum spread of the selected momentum states, we define $\hat{\psi}_{\uparrow,\downarrow}^{\dagger}$ and $\hat{\psi}_{\uparrow,\downarrow}$ as the operators for creating and annihilating an atom in momentum states ||$\ket{\uparrow}\equiv\ket{p_{0}+\hbar k}$ and ||$\ket{\downarrow}\equiv\ket{p_{0}-\hbar k}$. For mapping to a pseudo-spin model, we define ladder operators $\hat{J}_{+}=\hat{\psi}_{\uparrow}^{\dagger}\hat{\psi}_{\downarrow}$, $\hat{J}_{-}=\hat{\psi}_{\downarrow}^{\dagger}\hat{\psi}_{\uparrow}$ and spin projection operators $\hat{J}_{x}=\frac{1}{2}\left(\hat{J}_{+}+\hat{J}_{-}\right)$, $\hat{J}_{y}=\frac{1}{2i}\left(\hat{J}_{+}-\hat{J}_{-}\right)$ and $\hat{J}_{z}=\frac{1}{2}\left(\hat{\psi}_{\uparrow}^{\dagger}\hat{\psi}_{\uparrow}-\hat{\psi}_{\downarrow}^{\dagger}\hat{\psi}_{\downarrow}\right)$.
@@
As the atoms move along the cavity axis, the density grating is periodically aligned to the standing-wave of the cavity mode, leading to modulation of the cavity resonance frequency at the two-photon Doppler frequency $\omega_{z}=2kp_{0}/m\approx 2\pi\times 500~{}$kHz, where $m$ is the mass of ${}^{87}$b. To have this modulation to mediate an effective atom-atom interaction, we typically apply two $\sigma^{+}$ polarized dressing laser tones (see Fig. 1(A)) at frequencies $\omega_{1,2}$ within a few MHz of the cavity resonance frequency (see Fig. 1(B)), with complex amplitudes $\alpha_{1,2}$ corresponding to the field that would be established inside the cavity were no atoms in the cavity and implicit units $\sqrt{\mathrm{photons}}$. The atom–induced cavity frequency modulation leads to the generation of modulation sideband tones at frequencies $\omega_{1,2}\pm\omega_{z}.$ In the following simplifications, we will assume $\omega_{2}>\omega_{1}$.
The key insight is that different combinations of the dressing lasers and their atom-induced sideband tones will induce different virtual four-photon processes which will manifest as all-to-all exchange interactions $\hat{J}_{+}\hat{J}_{-}$ Thompson2023MomentumExchange and pair-raising $\hat{J}_{+}\hat{J}_{+}$ (lowering $\hat{J}_{-}\hat{J}_{-}$) processes as shown in Fig. 1(A,C). After adiabatically eliminating the cavity fields using second order perturbation theory (see Thompson2023MomentumExchange and Methods), we obtain an effective time-dependent atom-only Hamiltonian in an appropriate frame rotating at $\omega_{z}$
@@
Here, $\Delta_{\bar{c}}=\left(\omega_{2}+\omega_{1}\right)/2-\omega_{c}$ is the average detuning of the two dressing lasers from cavity resonance, typically set to be less than 1 MHz. $\delta=\left(\omega_{2}-\omega_{1}\right)-2\omega_{z}$ is the detuning from four-photon resonance, $g_{0}=2\pi\times 0.48$ MHz is the maximal atom-cavity Rabi coupling at an anti-node of the cavity mode, and $\phi_{\mathrm{int}}=\arg\left(\alpha_{2}\alpha_{1}^{*}\right)-\phi_{\mathrm{B}}$ is the differential phase between the two dressing laser tones relative to the phase of Bragg coupling $\phi_{\mathrm{B}}$ which forms initial density grating (see Methods). Collective cavity dissipation is dealt with separately (see Methods).
@@
where we defined the collective angular momentum operator, $\hat{\bf J}=\{\hat{J}_{x},\hat{J}_{y},\hat{J}_{z}\}$ and introduced the collective Heisenberg interaction $\hat{\bf J}\cdot\hat{\bf J}$. The latter acts as a constant for any eigenstate of the collective angular momentum operator $\hat{\bf J}\cdot\hat{\bf J}$, such as the collective states with eigenvalue $N/2(N/2+1)$. In the presence of single particle inhomogeneities, this term opens a many-body gap that help promote spin locking which we explored before for the protection of coherences against dephasing norcia2018cavity ; Thompson2023MomentumExchange . Here we instead focus on the fully collective dynamics and therefore without loss of generality we can add a generic $\chi_{z}\hat{\bf J}\cdot\hat{\bf J}$ term without affecting the dynamics. As such, in our system we are able to engineer dynamics governed by an XYZ Hamiltonian $\hat{H}=\chi_{x}\hat{J}^{2}_{x}+\chi_{y}\hat{J}^{2}_{y}+\chi_{z}\hat{J}^{2}_{z}$ with interaction strengths $\chi_{x}=\left(\chi_{e}+\chi_{p}+\chi_{z}\right)$ and $\chi_{y}=\left(\chi_{e}-\chi_{p}+\chi_{z}\right)$. The XYZ Hamiltonian is highly tunable by simply adjusting the relative power in the two applied dressing lasers since $\chi_{e}$ scales as $\left|\alpha_{2}\right|^{2}$, $\left|\alpha_{1}\right|^{2}$ and $\chi_{p}$ scales as $\left|\alpha_{2}\alpha_{1}\right|$.
@@
Mean-field dynamics

### v2

@@ Abstract
Quantum simulation using synthetic quantum systems offers unique opportunities to explore open questions in many-body physics and a path for the generation of useful entangled states. Nevertheless, so far many quantum simulators have been fundamentally limited in the models they can mimic. Here, we are able to realize an all-to-all interaction with arbitrary quadratic Hamiltonian or effectively an infinite range tunable Heisenberg XYZ model. This is accomplished by engineering cavity-mediated four-photon interactions between 700 rubidium atoms in which we harness a pair of momentum states as the effective pseudo spin or qubit degree of freedom. Using this capability we realize for the first time the so-called two-axis counter-twisting model at the mean-field level. The versatility of our platform to include more than two relevant momentum states, combined with the flexibility of the simulated Hamiltonians by adding cavity tones opens rich opportunities for quantum simulation and quantum sensing in matter-wave interferometers and other quantum sensors such as optical clocks and magnetometers.
thanks: These authors contributed equally.thanks: These authors contributed equally.
I Introduction
@@
The ability to create and control different many-body interactions is key for entanglement generation, optimization, quantum sensing, and quantum simulation. Long-range interacting systems have recently shown much promise for the engineering of Hamiltonians that generate interesting correlations that propagate across the system. Several experimental platforms are making rapid progress, including Rydberg atoms [1, 2, 3], polar molecules [4, 5], trapped ions [6, 7], cavity QED systems [8] and defect centers in solids [9, 10]. Short range contact interactions in ultra-cold atomic systems are another promising approach [11, 12, 13], however, it remains an open challenge to reach the sufficiently low-temperatures.
@@
So far, most efforts to engineer Hamiltonians have been limited to XXZ spin models or models that feature both exchange and Ising interactions. Common to these models is the fact that the total magnetization of the spin ensemble is preserved. However, limited progress has been achieved in engineering more general spin models, such as XYZ models, which can break both SU(2), and U(1) symmetries and lead to more general ground-state and out-of-equilibrium many-body behaviors. A few exceptions include experiments in Rydberg atoms using two-color dressing [14] with dynamics limited to pairs of atoms, or Floquet engineering in disordered arrays [2].
Here we experimentally show that photon-mediated interactions between atoms inside an optical cavity can realize a tunable all-to-all Heisenberg XYZ Hamiltonian, a model that offers rich avenues for the exploration of many-body physics and quantum metrology. We demonstrate the tunability of the XYZ Hamiltonian by explicitly mapping out the evolution of Bloch vectors on the Bloch sphere at the mean-field level for several control parameters that realize different canonical spin Hamiltonians.
@@
Collective XYZ Hamiltonians can in fact provide important gains in the context of quantum sensing and simulation applications. For instance, the so called two-axis counter-twisting (TACT) model features non-trivial and exponentially fast phase-space dynamics as compared to SU(2) and U(1) preserving models such as spin exchange or Ising interactions [15, 16, 17, 18]. The TACT model was originally proposed more than 30 years ago for generating squeezed states [19], as well as a simpler version, the so-called one-axis twisting (OAT) or collective Ising interaction [20, 19]. Only the TACT model is predicted to produce squeezed states approaching the Heisenberg limit. However, there has been no demonstration of genuine TACT in any platforms due to the challenge of how to realize this more complex non-linear Hamiltonian. Up to date only the OAT model has been realized in experiments including trapped ions [21, 22], Bose-Einstein Condensates [23, 24], atomic cavity-QED [15, 25, 17, 18], superconducting qubits [26] and optical interferometers [27]. Experiments have so far only approximated TACT dynamics locally on the Bloch sphere by combining OAT with a transverse drive (the so-called Lipkin-Meshkov-Glick (LMG) model) [28, 29, 30] or equivalent approximations in spin-nematic dynamics in higher spin systems [31, 32]. Here, as a particular case of the Heisenberg XYZ model, we are able to experimentally realize the TACT model using only two dressing lasers. This is a greatly simplified approach as compared to prior theory proposals in atom-cavity systems using four laser tones [33, 34].
@@
The pseudo-spin system here is novel, consisting of two momentum states of atoms freely-falling inside the cavity [35, 36], making our results of great interest for Bragg matter-wave interferometers [37] that are important for both inertial navigation and fundamental science such as searches for dark matter and dark energy, detection of gravitational waves, and determination of the fine structure constant [38, 39]. This approach can also be straightforwardly applied to systems with additional internal levels, making it ideal for developing next-generation quantum-enhanced sensors for technology and exploring a broad range of science from atomic clocks [40, 41] and magnetometers [42, 43, 44] to geodesy [45].
@@
Figure 1: Experimental overview. (A) Illustration of the microscopic momentum pair raising process described by the Hamiltonian $\hat{J}_{+}\hat{J}_{+}$, using two momentum states $p_{0}\pm\hbar k$ as a pseudo-spin 1/2 degree of freedom.Initially, two atoms $i$ and $j$ are in the same momentum state along the cavity axis. Dressing lasers are applied to the cavity (red and blue arrows) that allow atom $i$ to absorb a dressing laser photon and emit a photon (squiggly line) into the cavity such that the net photon recoil flips its momentum state by $2\hbar k$. The emitted photon is absorbed by atom $j$, also flipping its momentum state by $2\hbar k$. There also exists the separate momentum exchange process $\hat{J}_{+}\hat{J}_{-}$, where atoms initially in opposite momentum states flip their momentum states by emitting and absorbing photons explored in [46]. (B) Frequency diagram of the applied dressing lasers with frequencies $\omega_{1,2}$ and coherent state amplitudes $\alpha_{1,2}$. The emitted photons (squiggly line) are Doppler shifted by $\omega_{z}$ from the dressing laser frequencies and detuned by $\Delta_{c}$ from the cavity resonance frequency $\omega_{c}$. The cavity is far detuned from the atomic transition frequency $\omega_{a}$. (C) Representation of the emission/absorption processes described in (A) but depicted in terms of the atomic energy versus momentum for atoms $i$ and $j$. The reverse process is also allowed, giving rise to a collective lowering operator described by $\hat{J}_{-}\hat{J}_{-}$. (D) Dynamics induced by two-axis counter twisting in the form of a Hamiltonian $\hat{J}_{x}^{2}-\hat{J}_{z}^{2}$ is represented on a Bloch sphere with north and south poles defined by |$\ket{p_{0}\pm\hbar k}$. Top right shows the local circular flows around the stable point at $-\hat{x}$. The local flows for unstable saddle points at $\hat{y}$ with exponential squeezing and anti-squeezing along $\hat{x}\pm\hat{z}$ are shown in bottom right. (E) Observation of the four-photon resonance that generates pair raising and lowering processes. We scan the dressing laser frequency difference to vary the detuning $\delta$. With equal dressing laser amplitudes corresponding to realizing the $\hat{J}_{x}^{2}$ Hamiltonian when $\delta=0$, we see clear resonances in the observed change of the Bloch vector’s azimuthal angle $\phi$ for a Bloch vector prepared near the south pole ($\theta=\pi/4$, red data points) and near the north pole ($\theta=3\pi/4$, blue data points). All error bars reported are $1\sigma$ uncertainties. Simulation results are shown in solid lines with shaded area allowing for 5% uncertainty in interaction strength. See supplement for details. For all the Hamiltonians engineered later, we will focus on the resonant case with $\delta=0$ as highlighted by the red solid line
II Experimental setup
@@
In the experiment, 87Rb atoms are laser-cooled inside a vertically-oriented two-mirror standing wave cavity, see Fig. 1(A) and [46, 35]. A repulsive intra-cavity doughnut dipole trap confines the atoms radially, but allows the atoms to fall along the cavity axis. To prepare atoms in a well-defined momentum state, a pair of laser beams are injected into the cavity to drive velocity-dependent two-photon Raman transitions between ground hyperfine states |$\ket{F=1,m_{F}=0}$ and |$\ket{F=2,m_{F}=0}$. After removing the unselected atoms with a resonant laser push beam, successive microwave pulses are applied to prepare the internal states of the selected atoms in the ground hyperfine state |$\ket{F=2,m_{F}=2}$ (see supplement and [46]).
@@
With about 700 atoms centered at momentum $p_{0}-\hbar k$, another pair of laser beams is injected along the cavity axis to drive two-photon Bragg transitions connecting the two momentum states |$\ket{p_{0}\pm\hbar k}$ which defines a two-level spin-1/2 system [35, 46]. Here the average momentum is $p_{0}$, $\hbar$ is the reduced Plank constant, and $k=2\pi/\lambda$ where $\lambda$ is the wavelength of the Bragg laser beams. Due to finite momentum spread, one can consider momentum wave packets centered at these two momentum states as considered in detail previously [46]. Ignoring the finite momentum spread of the selected momentum states, we define $\hat{\psi}_{\uparrow,\downarrow}^{\dagger}$ and $\hat{\psi}_{\uparrow,\downarrow}$ as the operators for creating and annihilating an atom in momentum states ||$\ket{\uparrow}\equiv\ket{p_{0}+\hbar k}$ and ||$\ket{\downarrow}\equiv\ket{p_{0}-\hbar k}$. For mapping to a pseudo-spin model, we define ladder operators $\hat{J}_{+}=\hat{\psi}_{\uparrow}^{\dagger}\hat{\psi}_{\downarrow}$, $\hat{J}_{-}=\hat{\psi}_{\downarrow}^{\dagger}\hat{\psi}_{\uparrow}$ and spin projection operators $\hat{J}_{x}=\frac{1}{2}\left(\hat{J}_{+}+\hat{J}_{-}\right)$, $\hat{J}_{y}=\frac{1}{2i}\left(\hat{J}_{+}-\hat{J}_{-}\right)$ and $\hat{J}_{z}=\frac{1}{2}\left(\hat{\psi}_{\uparrow}^{\dagger}\hat{\psi}_{\uparrow}-\hat{\psi}_{\downarrow}^{\dagger}\hat{\psi}_{\downarrow}\right)$.
@@
As the atoms move along the cavity axis, the density grating is periodically aligned to the standing-wave of the cavity mode, leading to modulation of the cavity resonance frequency at the two-photon Doppler frequency $\omega_{z}=2kp_{0}/m\approx 2\pi\times 500~{}$kHz, where $m$ is the mass of 87Rb. To have this modulation to mediate an effective atom-atom interaction, we typically apply two $\sigma^{+}$ polarized dressing laser tones (see Fig. 1(A)) at frequencies $\omega_{1,2}$ (see Fig. 1(B)), with complex amplitudes $\alpha_{1,2}$ corresponding to the field that would be established inside the cavity were no atoms in the cavity and implicit units $\sqrt{\mathrm{photons}}$. The atom-induced cavity frequency modulation leads to the generation of modulation sideband tones at frequencies $\omega_{1,2}\pm\omega_{z}.$ In the following simplifications, we will assume $\omega_{2}>\omega_{1}$.
The key insight is that different combinations of the dressing lasers and their atom-induced sideband tones will induce different virtual four-photon processes which will manifest as all-to-all exchange interactions $\hat{J}_{+}\hat{J}_{-}$ [46] and pair-raising $\hat{J}_{+}\hat{J}_{+}$ (lowering $\hat{J}_{-}\hat{J}_{-}$) processes as shown in Fig. 1(A,C). After adiabatically eliminating the cavity fields using second order perturbation theory (see [46] and supplement), we obtain an effective time-dependent atom-only Hamiltonian in an appropriate frame rotating at $\omega_{z}$
@@
Here, $\Delta_{\bar{c}}=\left(\omega_{2}+\omega_{1}\right)/2-\omega_{c}$ is the average detuning of the two dressing lasers from cavity resonance, typically set to be less than 1 MHz. $\delta=\left(\omega_{2}-\omega_{1}\right)-2\omega_{z}$ is the detuning from four-photon resonance, $g_{0}=2\pi\times 0.48$ MHz is the maximal atom-cavity Rabi coupling at an anti-node of the cavity mode, and $\phi_{\mathrm{int}}=\arg\left(\alpha_{2}\alpha_{1}^{*}\right)-\phi_{\mathrm{B}}$ is the differential phase between the two dressing laser tones relative to the phase of Bragg coupling $\phi_{\mathrm{B}}$ which forms initial density grating (see supplement). Collective cavity dissipation is dealt with separately (see supplement).
@@
where we defined the collective angular momentum operator, $\hat{\bf J}=\{\hat{J}_{x},\hat{J}_{y},\hat{J}_{z}\}$ and introduced the collective Heisenberg interaction $\hat{\bf J}\cdot\hat{\bf J}$. The latter acts as a constant for any eigenstate of the collective angular momentum operator $\hat{\bf J}\cdot\hat{\bf J}$, such as the collective states with eigenvalue $N/2(N/2+1)$. In the presence of single particle inhomogeneities, this term opens a many-body gap that help promote spin locking which we explored before for the protection of coherences against dephasing [17, 46, 18]. Here we instead focus on the fully collective dynamics and therefore without loss of generality we can add a generic $\chi_{z}\hat{\bf J}\cdot\hat{\bf J}$ term without affecting the dynamics. As such, in our system we are able to engineer dynamics governed by an XYZ Hamiltonian $\hat{H}=\chi_{x}\hat{J}^{2}_{x}+\chi_{y}\hat{J}^{2}_{y}+\chi_{z}\hat{J}^{2}_{z}$ with interaction strengths $\chi_{x}=\left(\chi_{e}+\chi_{p}+\chi_{z}\right)$ and $\chi_{y}=\left(\chi_{e}-\chi_{p}+\chi_{z}\right)$. The XYZ Hamiltonian is highly tunable by simply adjusting the relative power in the two applied dressing lasers since $\chi_{e}$ scales as $\left|\alpha_{2}\right|^{2}$, $\left|\alpha_{1}\right|^{2}$ and $\chi_{p}$ scales as $\left|\alpha_{2}\alpha_{1}\right|$.
To realize the XYZ Hamiltonians of Fig. 2 and 3, we apply the two dressing laser tones separated by $\Delta_{p}=\omega_{2}-\omega_{1}=2\omega_{z}\approx 2\pi\times 1$ MHz, and with average dressing laser detuning from the cavity set to $\Delta_{\bar{c}}=2\pi\times 200$ kHz as shown in Fig. 1. Because the atoms are accelerating due to gravity, $\omega_{z}$ changes linearly in time at a rate $d\omega_{z}/dt=2\pi\times 25.11~{}$kHz/ms. To compensate for this, we linearly ramp the dressing laser separation at a rate $d\Delta_{p}/dt=2\,d\omega_{z}/dt$. Similarly, the difference frequency of the applied Bragg coupling tones is ramped at $d\omega_{z}/dt$. For the TACT Hamiltonian dynamics presented in Fig. 2 and 3, the total power of the incident dressing lasers was approximately 200 pW or 900 photons/$\mu$s to realize an interaction strength of $\chi_{p}=2\pi\times 1.25$ Hz.
@@
III Mean-field dynamics